# CICLe: Conformal In-Context Learning for Largescale Multi-Class Food Risk Classification

### v2

@@ Abstract
Contaminated or adulterated food poses a substantial risk to human health. Given sets of labeled web texts for training, Machine Learning and Natural Language Processing can be applied to automatically detect such risks. We publish a dataset of 7,546 short texts describing public food recall announcements. Each text is manually labeled, on two granularity levels (coarse and fine), for food products and hazards that the recall corresponds to. We describe the dataset and benchmark naive, traditional, and Transformer models. Based on our analysis, Logistic Regression based on a tf-idf representation outperforms RoBERTa and XLM-R on classes with low support. Finally, we discuss different prompting strategies and present an LLM-in-the-loop framework, based on Conformal Prediction, which boosts the performance of the base classifier while reducing energy consumption compared to normal prompting.
@@
Korbinian Randl,1 John Pavlopoulos,2,1,3 Aron Henriksson,1 Tony Lindgren1 1 Stockholm University, Borgarfjordsgatan 12, 164 07 Kista, Sweden {korbinian.randl,ioannis,aronhen,tony}@dsv.su.se 2Athens University of Economics and Business, Patission 76, Athens 104 34, Greece annis@aueb.gr 3Archimedes/Athena RC
@@ 1 Introduction
Food-bourne illnesses and contaminated food pose a serious threat to human health and lead to thousands of deaths Majowicz et al. (2014); de Noordhout et al. (2014). Natural Language Processing (NLP) solutions based on Machine and Deep Learning (ML, DL) or Large Language Models (LLMs) enable fast responses to new threats by generating warnings from publicly available texts on the internet Maharana et al. (2019); Tao et al. (2023). These texts, however, as we show, can be noisy and are characterized by thousands of classes. While we are aware of existing text datasets on the topic of food-bourne illnesses Hu et al. (2022), they focus only on the detection of such illnesses and not on their categorization. Categorization is the natural next step after detection and is needed to properly address food risks as different kinds of hazards may warrant different actions. In this work, we address this classification task and provide the following main contributions:
@@
    We discuss different approaches to few-shot prompting with GPT-3.5 on our data and propose reducing the number of relevant classes for few-shot prompting using Conformal Prediction (Vovk et al., 2005, CP). We show that this enables resource-friendly prompting at comparably high $F_{1}$ by boosting a base classifier’s performance in a LLM-in-the-loop framework.
@@ 2 Background and Related Work
has traditionally focused on tabular and image data Zhou et al. (2019); Jin et al. (2020); Wang et al. (2022), where food risk classification is usually presented as a detection task: rather than identifying specific hazards, the task is formulated as the binary classification of the presence or absence of an incident Maharana et al. (2019); Tao et al. (2023); Wang et al. (2023). Focusing on the text modality, we found literature exploring such detection tasks based on Amazon reviews matched with FDA food recall announcements Maharana et al. (2019) as well as Twitter data, labeled using instances from the US National Outbreak Reporting System Tao et al. (2023). In both of these papers, the data is highly biased towards the negative class (i.e. no incident). Furthermore, the description of the method for labeling the Twitter texts in Tao et al. (2023) lacks important detail, as it is unclear on what grounds and how texts and events are matched. We only found one study on the actual classification of hazards: Xiong et al. (2023) which use a hierarchical Transformer model to classify Chinese news texts into four categories of hazards. The authors do not share further details on their data, which makes an assessment of the paper’s validity hard. Overall, the field of food risk classification based on text lacks reproducible work on fine-grained prediction of hazards based on openly accessible data.
@@
Recently, LLMs such as Llama 2 Touvron et al. (2023), PaLM Chowdhery et al. (2022), and GPT-family models Brown et al. (2020) have been shown to exceed the capabilities of smaller Transformers even without further fine-tuning. That is, by simply providing a context of a few labeled samples per class, LLMs can predict the classes of unseen samples within this context. As LLMs are usually text-to-text Transformers, the context is provided directly in each prompt. This paradigm is commonly referred to as few-shot prompting or in-context learning Brown et al. (2020). Since the detection of LLMs’ few-shot capabilities, newer LLMs are often designed for high few-shot performance Gao et al. (2021); Chowdhery et al. (2022).
Current work on few-shot prompt engineering
focuses mostly on creating/finding the optimal few-shot samples from the training data. Ahmed et al. (2023) proposed a workflow for automatically finding similar samples from the pool of labeled “training" samples for code summarization. Shi et al. (2023) focused on automatic generation of Chain of Thought (CoT) labels for samples in reasoning tasks. CoT is another prompting paradigm that asks the LLM to provide a chain of reasoning before delivering the prediction and has been shown to drastically improve reasoning performance Wei et al. (2022). Nevertheless, to the best of our knowledge, there is no previous work on how to leverage few-shot prompting in LLMs for multi-class prediction problems with a high number of classes. While prompting with long contexts may “confuse” the LLM (Liu et al., 2023), the context window may even be too small to support a view of all classes in such cases. Furthermore, the previous approaches focus on selecting specific samples for the prompt rather than relevant classes for the prompt.
@@ 3.1 Data Description
hazard: A fine-grained description of the hazards mentioned in the texts comprising $261$ classes. hazard-category: A categorized version of the hazard classification task comprising $10$ classes. product: A fine-grained description of the products mentioned in the texts comprising 1,256 classes. product-category: A categorized version of the product classification task comprising $22$ classes. The columns hazard-title and product-title comprise character spans, generated based on feature importance of a Logistic Regression (LR) classifier (see §3.1.1). These signify parts of the title that are important for the hazard and product classification. As the fine-grained tasks, hazard and product, have very low support for many classes, they may require further pre-processing (e.g. clustering/filtering of labels) dependent on the application. The dataset, publicly released under a Creative Commons BY-NC-SA 4.0 license, comprises also metadata, such as the release date of the text (columns year, month, and day), the language of the text (column language), and the country of issue (column country). Appendix A contains more detailed statistics, as well as sample texts.
@@ 3.1.1 Ground Truth
Fine-grained labels were assigned per web-domain by a domain expert of Agroknow and then grouped to the final labels using internal ontologies. [^4] Random checks of the labels were performed by more experienced curators for quality assurance. In order to measure inter-annotator agreement (IAA), we asked a senior curator involved in the labeling process and one independent expert to label a sample of 30 texts (stratified over all labels) independently of each other. Comparing the annotator’s assessment to the labels in the dataset shows moderate agreement (Cohen’s $\kappa$: 0.76 on hazard-category, 0.63 on product-category). The expert’s assessment is intuitively slightly less aligned (Cohen’s $\kappa$: 0.63 on hazard-category, 0.60 on product-category) but still acceptable. Expert and Annotator show a $\kappa$ of 0.59 on hazard-category, 0.77 on product-category. The raw data of our IAA study are available along with the dataset.
@@ 4.1 The Task
The basic problem we address in this paper is extreme multi-class classification on heavily imbalanced data. More formally, given a number of training-texts $T_{i},~{}i\in\{1,2,\dots,N\}$ and their corresponding classes $y_{i}$, we aim to train a classifier $\mathrm{f}(T_{i},\cdot)=\hat{y}_{i}$ that minimizes the error $|y_{i}-\hat{y}_{i}|$. For standard ML classifiers, the function $\mathrm{f}(T_{i},\cdot)$ is usually a two-step process with the first step mapping the text to a machine-readable embedding vector $X_{i}=\mathrm{e}(T_{i})$, and the second one involving the learning process on $X_{i}$: $\mathrm{f}(T_{i},\cdot)=\mathrm{f^{\prime}}(X_{i},\cdot)$. Both the class labels $y_{i}$ and predictions $\hat{y}_{i}$ are integer scalars in $\{1,2,\dots,M\}$. In all our classification tasks, we have at least $M\geq 10$.
4.2 ML Classifiers
As a naive baseline, we report the performance of two classifiers. The random classifier (RANDOM) yields a random integer in $\{1,2,\dots,M\}$ for each $\hat{y}_{i}$. The majority baseline (MAJORITY), chooses each $\hat{y}_{i}$ to be the class with the highest number of samples in the training data.
@@ 4.2.1 Traditional ML Classifiers
We use Bag-of-Words (BOW) and Term Frequency - Inverse Document Frequency (Spärck Jones, 1972, TF-IDF) encodings (see Appendix C), combined with $k$-Nearest Neighbours (KNN), LR, or Support Vector Machine (SVM) classifiers. [^5] For the classifiers, we use the implementation from the Python library scikit-learn Pedregosa et al. (2011). For KNN we optimize over $k\in\{2,4,8\}$ neighbours based on cosine similarity of the embedding vectors. For LR, we use a ‘liblinear’ solver and optimize $C\in\{0.5,1.0,2.0\}$ for both L1 and L2 regularization. For the SVMs, we use a linear kernel and optimize the parameter $C\in\{0.5,1.0,2.0\}$ for L2 regularization on the validation splits (defined in §5.1). For LR and SVM, we use a one-vs-all approach to classification, which means that we train one binary classifier for each class in the predicted classification task and choose the class with the highest probability.
@@ 4.3 Prompting
We also employ few-shot prompting with gpt-3.5-turbo-instruct. [^7] Accessed on 02-2024 through OpenAI’s python API. For all our prompting examples, we use a temperature value of 00. The naive approach is to provide a context describing the classification task followed by the two most similar text-label pairs (based on cosine similarity of the TF-IDF-embeddings) per class from the training data. These examples are ordered from the most similar to least similar in the prompt (dubbed GPT-ALL).
@@
GPT-ALL is applicable only to coarse tasks due to the size of the context window: for the hazard task, we would have at most $\frac{4,096~{}\mathrm{tokens}}{261~{}\mathrm{classes}}=15.7$ tokens per class which is insufficient for providing two example texts. Furthermore, reduced prompt sizes can lead to increased energy efficiency, as the size of the attention weights computed in the Transformer is input size $\times$ output size (Vaswani et al., 2017). Lastly, meaningfully reducing the classes in the prompt can increase predictive performance, as the probability of choosing the right class increases (Appendix D). Therefore, we use three different kinds of prompts (see Figure 3) with the third kind (GPT-CICLe) being novel. GPT-SIM-$k$ is simply the GPT-ALL prompt shortened by only keeping the $k$ most similar few-shot-samples. GPT-MAX-$k$ selects the two most similar few-shot-samples from the $k$ most probable class labels as predicted by a base classifier. The few-shot samples are ordered from most to least probable, which yields $2\cdot k$ few-shot samples per prompt.
@@ 4.3.1 Our proposed GPT-CICLe
We propose using Conformal Prediction (CP) Vovk et al. (2005) to adapt the context length to a base classifier’s certainty on the predicted classes. CP uses the concept of a non-conformity measure $\delta(y_{i},\hat{y}_{i})$ to create sets of predicted classes, which contain the true class with a probability of $p\approx 1-\alpha$.
In our case, this non-conformity measure is simply the base classifier’s uncertainty on the true class: $\delta(y_{i},\hat{y}_{i})=1-P(\hat{y}_{i}=y_{i})$. Here, $\hat{y}_{i}$ are the predictions and $y_{i}$ are the corresponding true labels. If one defines $q=\frac{\lceil{(N+1)(1-\alpha)}\rceil}{N}$, and $\hat{q}$ as the $q^{th}$ empirical quantile of $\{\delta(Y_{i},\hat{Y}_{i})|1\leq i\leq N\}$, it can be shown that for any prediction $\hat{Y}$ of the classifier on an unknown sample, a set of classes $\{j|\hat{y}_{j}\geq 1-\hat{q}\}$ contains the true class with probability $p\geq 1-\alpha$ Vovk et al. (2005). This means that we accept an upper bound of $1-\alpha$ on prompting accuracy, which we trade for adaptive context length based on sample difficulty.
@@ 5.1 Experimental Setup
For training and evaluation of our ML models, we apply 5-fold cross-validation (CV) to create 5 train-test splits. From each of these 5 training sets, we create a validation set using $10\%$ holdout. To reduce the computational cost, we only test our prompting methods other than GPT-ALL on the first CV split, which corresponds to a $20\%$ holdout. We use stratification on the *-category classification tasks as those provide a sufficient number of samples for splitting in each class. We use the same splits for hazard-category and hazard, as well as product-category and product to keep the results comparable. Our classifiers provide baseline performance on the dataset. In order to demonstrate the effect of class imbalance on performance, we do not employ balancing methods like oversampling or class weights during training.

### v3

@@ Abstract
Contaminated or adulterated food poses a substantial risk to human health. Given sets of labeled web texts for training, Machine Learning and Natural Language Processing can be applied to automatically detect such risks. We publish a dataset of 7,546 short texts describing public food recall announcements. Each text is manually labeled, on two granularity levels (coarse and fine), for food products and hazards that the recall corresponds to. We describe the dataset and benchmark naive, traditional, and Transformer models. Based on our analysis, Logistic Regression with a tf-idf representation outperforms RoBERTa and XLM-R on classes with low support. Finally, we discuss different prompting strategies and present an LLM-in-the-loop framework, based on Conformal Prediction, which boosts the performance of the base classifier while reducing energy consumption compared to normal prompting.
@@
Korbinian Randl,1 John Pavlopoulos,1,2,3 Aron Henriksson,1 Tony Lindgren1 1 Stockholm University, Borgarfjordsgatan 12, 164 07 Kista, Sweden {korbinian.randl,ioannis,aronhen,tony}@dsv.su.se 2Athens University of Economics and Business, Patission 76, Athens 104 34, Greece 3Archimedes/Athena RC
@@ 1 Introduction
Food-bourne illnesses and contaminated food pose a serious threat to human health and lead to thousands of deaths Majowicz et al. (2014); de Noordhout et al. (2014). Natural Language Processing (NLP) solutions based on Machine and Deep Learning (ML, DL) or Large Language Models (LLMs) enable fast responses to new threats by generating warnings from publicly available texts on the internet Maharana et al. (2019); Tao et al. (2023). These texts, however, as we show, can be noisy and are characterized by thousands of classes. Categorization of incidents into these classes is needed to properly address food risks, as different food risks require different countermeasures with different urgency. For example, while a fraudulent import of foodstuffs usually imposes a very low risk to consumer health, Listeria or Salmonella contamination is a serious risk to consumer health and can even lead to death. Automatic classification of the hazard involved in food recalls can help estimate the severity of food risks and lead to faster and more targeted measures. While we are aware of existing text datasets on the topic of food-bourne illnesses Hu et al. (2022), they focus only on the detection of such illnesses and not on their categorization. Publicly available official food recall announcements, and specifically their shorter titles, contain detailed information on food risks and the specific products involved and may be used as a basis for data-driven ML approaches. Acknowledging this research gap, this paper is specifically addressing the following research questions:
R1: Are titles of food recall articles sufficient for classification of the food hazards and food products involved?
  • R1.1: Do the titles of official announcements of food hazards comprise enough information for their classification?
  • R1.2: What is the performance of ML and DL classification models trained on the presented dataset?
As LLMs like GPT-4, and Llama Touvron et al. (2023) represent the current state-of-the-art in NLP, but are also known for their high resource consumption, we further explore the following question:
R2: Can LLMs be used to address or improve automatic food hazard classification?
  • R2.1: What is the performance of LLMs for the classification tasks?
  • R2.2: How can ML and DL be used to increase resource efficiency without harming the accuracy of LLMs for the task?
In this work, we address the above task of classifying food hazards and the involved products from food recall texts and provide the following main contributions:
@@
    We discuss different approaches to few-shot prompting with GPT-3.5 on our data and propose reducing the number of relevant classes for few-shot prompting using Conformal Prediction (CP). We show that this enables resource-friendly prompting at comparably high $F_{1}$ by boosting a base classifier’s performance in a LLM-in-the-loop framework.
@@ 2 Background and Related Work
has traditionally focused on tabular and image data Zhou et al. (2019); Jin et al. (2020); Wang et al. (2022), where food risk classification is usually presented as a detection task: rather than identifying specific hazards, the task is formulated as the binary classification of the presence or absence of an incident Maharana et al. (2019); Tao et al. (2023); Wang et al. (2023). Focusing on the text modality, we found literature exploring such detection tasks based on Amazon reviews matched with FDA food recall announcements Maharana et al. (2019), as well as Twitter data, labeled using instances from the US National Outbreak Reporting System Tao et al. (2023). In both of these papers, the data is highly unbalanced towards the negative class (i.e. no incident). Furthermore, the description of the method for labeling the Twitter texts in Tao et al. (2023) lacks important details, as it is unclear on what grounds texts and events are matched. We only found one study on the actual classification of hazards: Xiong et al. (2023) which uses a hierarchical Transformer model to classify Chinese news texts into four categories of hazards. The authors do not share further details on their data, which makes an assessment of the paper’s validity hard. Overall, the field of food risk classification based on text lacks reproducible work on fine-grained prediction of hazards based on openly accessible data.
@@
Recently, LLMs such as Llama 2 Touvron et al. (2023), PaLM Chowdhery et al. (2022), and GPT-family models Brown et al. (2020) have been shown to exceed the capabilities of smaller Transformers even without further fine-tuning. That is, by simply providing a context of a few labeled samples per class, LLMs can predict the classes of unseen samples within this context. As LLMs are usually text-to-text Transformers, the context is provided directly in each prompt. This paradigm is commonly referred to as few-shot prompting or in-context learning Brown et al. (2020). Since the discovery of the few-shot capabilities of LLMs, newer LLMs are often designed for high few-shot performance Gao et al. (2021); Chowdhery et al. (2022).
Current work on few-shot prompting
focuses mostly on creating/finding the optimal few-shot samples from the training data. Ahmed et al. (2023) proposed a workflow for automatically finding similar samples from the pool of labeled “training” samples for code summarization. Shi et al. (2023) focused on automatic generation of Chain of Thought (CoT) labels for samples in reasoning tasks. CoT is another prompting paradigm that asks the LLM to provide a chain of reasoning before delivering the prediction and has been shown to drastically improve reasoning performance Wei et al. (2022). Nevertheless, to the best of our knowledge, there is no previous work on how to leverage few-shot prompting in LLMs for multi-class prediction problems with a high number of classes. While prompting with long contexts may “confuse” the LLM (Liu et al., 2023), the context window may even be too small to support a view of all classes in such cases. Furthermore, the previous approaches focus on selecting specific samples rather than relevant classes for the prompt.
The CP framework
Vovk et al. (2005) is a method for associating predictions of a classification algorithm with confidence guarantees. Essentially, CP selects sets of classes from the predicted labels of a classifier or regressor that statistically contain the true label with a predefined probability. We note, however, that the standard CP setting provides guarantees for the individual class labels; i.e., errors for one of the labels can, hence, exceed the set guarantees. A formulation, called Mondrian of CP, can make the guarantees hold also for the individual labels. Note that CP does a very good job of keeping the output (set of labels) small, while the prediction guarantees hold. It can be applied to any classification (or regression) model, as long as a calibration set is set aside together with a measure of the model’s prediction error with regard to the true labels. This measure is referred to as the “non-conformity function”. In our context, the stronger the model, and the better the non-conformity function, the fewer labels (hence, shots) will be produced with the same guarantee. We point to the work of Vovk et al. (2005); johansson2014regression; 8260735 for more information on CP.
@@ 3.1 Data Description
hazard: A fine-grained description of the hazards mentioned in the texts comprising $261$ classes. hazard-category: A categorized version of the hazard classification task comprising $10$ classes. product: A fine-grained description of the products mentioned in the texts comprising 1,256 classes. product-category: A categorized version of the product classification task comprising $22$ classes. The columns hazard-title and product-title comprise character spans, generated based on feature importance of a Logistic Regression (LR) classifier (see §3.1.1). These signify parts of the title that are important for the hazard and product classification. As the fine-grained tasks, hazard and product, have very low support for many classes, they may require further pre-processing (e.g. clustering/filtering of labels) depending on the application. The dataset, publicly released under a Creative Commons BY-NC-SA 4.0 license, comprises also metadata, such as the release date of the text (columns year, month, and day), the language of the text (column language), and the country of issue (column country). Appendix A contains more detailed statistics, as well as sample texts.
@@ 3.1.1 Ground Truth
Fine-grained labels were assigned per web-domain by a domain expert of Agroknow and then grouped to the final labels using internal ontologies. [^4] Random checks of the labels were performed by more experienced curators for quality assurance. In order to measure inter-annotator agreement (IAA), we asked a senior curator involved in the labeling process and one independent expert to label a sample of 30 texts (stratified over both *-category labels) independently of each other. Comparing the annotator’s assessment to the labels in the dataset shows moderate agreement (Cohen’s $\kappa$: 0.76 on hazard-category, 0.63 on product-category). The expert’s assessment is intuitively slightly less aligned (Cohen’s $\kappa$: 0.63 on hazard-category, 0.60 on product-category) but still acceptable. Expert and Annotator show a $\kappa$ of 0.59 on hazard-category, 0.77 on product-category. The raw data of our IAA study are available along with the code.
@@ 4.1 The Task
The basic problem we address in this paper is extreme multi-class classification on heavily imbalanced data. More formally, given a number of training texts $T_{i},~{}i\in\{1,2,\dots,N\}$ and their corresponding classes $y_{i}$, we aim to train a classifier $\mathrm{f}(T_{i},\cdot)=\hat{y}_{i}$ that minimizes the error $|y_{i}-\hat{y}_{i}|$. For standard ML classifiers, the function $\mathrm{f}(T_{i},\cdot)$ is usually a two-step process with the first step mapping the text to a machine-readable embedding vector $X_{i}=\mathrm{e}(T_{i})$, and the second one involving the learning process on $X_{i}$: $\mathrm{f}(T_{i},\cdot)=\mathrm{f^{\prime}}(X_{i},\cdot)$. Both the class labels $y_{i}$ and predictions $\hat{y}_{i}$ are integer scalars in $\{1,2,\dots,M\}$. In all our classification tasks, we have at least $M\geq 10$.
4.2 Classifiers
As a naive baseline, we report the performance of two classifiers. The random classifier (RANDOM) yields a random integer in $\{1,2,\dots,M\}$ for each $\hat{y}_{i}$. The majority baseline (MAJORITY) chooses each $\hat{y}_{i}$ to be the class with the highest number of samples in the training data.
@@ 4.2.1 Traditional ML Classifiers
We use Bag-of-Words (BOW) and Term Frequency - Inverse Document Frequency (Spärck Jones, 1972, TF-IDF) encodings (see Appendix C), combined with $k$-Nearest Neighbours (KNN), LR, or Support Vector Machine (SVM) classifiers. [^5] For the classifiers, we use the implementation from the Python library scikit-learn Pedregosa et al. (2011). For KNN, we optimize over $k\in\{2,4,8\}$ neighbours based on cosine similarity of the embedding vectors. For LR, we use a ‘liblinear’ solver and optimize $C\in\{0.5,1.0,2.0\}$ for both L1 and L2 regularization. For the SVMs, we use a linear kernel and optimize the parameter $C\in\{0.5,1.0,2.0\}$ for L2 regularization on the validation splits (defined in §5.1). For LR and SVM, we use a one-vs-all approach to classification, which means that we train one binary classifier for each class in the predicted classification task and choose the class with the highest probability.
@@ 4.3 Prompting
Prompt engineering can lead to accurate classification results with limited training data. We investigate this path for our task by employing few-shot in-context learning with gpt-3.5-turbo-instruct. [^7] Accessed on 02-2024 through OpenAI’s Python API. For all our prompting examples, we use a temperature value of 00. The naive approach is to provide a context describing the classification task followed by the two most similar text-label pairs (based on cosine similarity of the TF-IDF-embeddings) per class from the training data. These examples are ordered from the most similar to least similar in the prompt (dubbed GPT-ALL).
@@
GPT-ALL is applicable only to coarse tasks due to the size of the context window: for the hazard task, we would have at most $\frac{4,096~{}\mathrm{tokens}}{261~{}\mathrm{classes}}=15.7$ tokens per class which is insufficient for providing two example texts. Furthermore, reduced prompt sizes can lead to increased energy efficiency, as the size of the attention weights computed in the Transformer is input size $\times$ output size (Vaswani et al., 2017). Lastly, meaningfully reducing the classes in the prompt can increase predictive performance, as the probability of choosing the right class increases (Appendix D). Therefore, we use three different kinds of prompts (see Figure 3) with the third kind (GPT-CICLe) being novel. GPT-SIM-$k$ is simply the GPT-ALL prompt shortened by only keeping the $k$ most similar few-shot samples. GPT-MAX-$k$ selects the two most similar few-shot samples from the $k$ most probable class labels as predicted by a base classifier. The few-shot samples are ordered from most to least probable, which yields $2\cdot k$ few-shot samples per prompt.
@@ 4.3.1 Our proposed GPT-CICLe
We consider that the better the few shots the LLM is provided with, the better its performance will be. Therefore, in this work, we propose using Conformal Prediction (CP) Vovk et al. (2005) to adapt the context length to a base classifier’s certainty on the predicted classes. CP uses the concept of a non-conformity measure $\delta(y_{i},\hat{y}_{i})$ to create sets of predicted classes, which contain the true class with a probability of $p\approx 1-\alpha$.
In our case, this non-conformity measure is simply the base classifier’s uncertainty on the true class:
$\delta(y_{i},\hat{y}_{i})=1-P(\hat{y}_{i}=y_{i}).$
Here, $\hat{y}_{i}$ are the predictions and $y_{i}$ are the corresponding true labels. If one defines
$q=\frac{\lceil{(N+1)(1-\alpha)}\rceil}{N},$
and $\hat{q}$ as the $q^{th}$ empirical quantile of $\{\delta(Y_{i},\hat{Y}_{i})|1\leq i\leq N\}$, it can be shown that for any prediction $\hat{Y}$ of the classifier on an unknown sample, a set of classes $\{j|\hat{y}_{j}\geq 1-\hat{q}\}$ contains the true class with probability $p\geq 1-\alpha$ Vovk et al. (2005). This means that we accept an upper bound of $1-\alpha$ on prompting accuracy, which we trade for adaptive context length based on sample difficulty.
@@ 5.1 Experimental Setup
For training and evaluation of our ML models, we apply 5-fold cross-validation (CV) to create 5 train-test splits. From each of these 5 training sets, we create a validation set using $10\%$ holdout which we use for hyperparameter tuning as described in Appendix C. To reduce the computational cost, we only test our prompting methods other than GPT-ALL on the first CV split, which corresponds to a $20\%$ holdout. We use stratification on the *-category classification tasks as those provide a sufficient number of samples for splitting in each class. We use the same splits for hazard-category and hazard, as well as product-category and product to keep the results comparable. Our classifiers provide baseline performance on the dataset. In order to demonstrate the effect of class imbalance on performance, we do not employ balancing methods like oversampling or class weights during training.